OpenStudy (anonymous):

integral x/(4th root (x+2) dx

OpenStudy (anonymous):

\[\int\limits_{}^{}x/\sqrt[4]{x+2}dx\]

OpenStudy (anonymous):

You can use a substitution: \[u = x+2\] \[x = u-2\] \[dx = du\] \[\int\limits_{}^{}(u-2)(u^{-1/4}) du\] Distribute, then integrate like you normally would When you finish integrating (don't forget your C's!), substitute \[u = x+2\] back in.